# Towards a Formal Foundation for Blockchain ZK Rollups

### v2

@@ 5 Design-time Testing with Alloy Models
There could be two issues arising from merely enforcing the upgrade after a set period of time. For the first example, consider the case where we are using a forced queue and upgradeability as described in Section 4.5. However, we deploy the upgrade based solely on the timeout (i.e., not taking care of locking and emptying the forced queue before deploying the upgrade). Figure 3 demonstrates the issue that might arise in that case produced by Alloy. Specifically, what happens is that first, we start with an empty forced queue, and then a user adds an input to the queue (Input_1). Next, an announcement occurs, and then the timeout of that announcement is reached. This enforces the upgrade, updating the blacklist to include Input_1. The result is that L2 becomes frozen because the head (Input_1) of the queue is blacklisted. Another problem could occur if the upgrade involves changing the functionality of the rollup and a user might submit a transaction to be forced before the upgrade because they disagree with the change and want to exit before the upgrade. However, because other transactions are in front of it (whether benignly or maliciously added), their transaction will not be processed. Additionally, the L2 might censor the user’s transactions, preventing the user from exiting before the upgrade.
These examples demonstrate severe issues with existing mechanisms. Our model presented in Section 4 is immune to the described attacks.
6 Discussion
Why now. The L2Beat Risk page indicates that the most widely used optimistic rollups already support mechanisms like the Forced Queue, referred to as “Sequencer failure” on L2Beat. Additionally, as we demonstrated in Section 4.5, the state-of-the-art upgradeability mechanisms can be exploited by malicious actors. Most rollups struggle to design an upgradeability mechanism that is both secure and compatible with other mechanisms. Our work provides a framework to help design and evaluate such mechanisms, along with a secure design. This is crucial because for rollups to truly inherit the security and censorship resistance of the L1 they are based on, they need secure implementation of the mechanisms discussed in Section 4. We hope that our work will encourage L2Beat to integrate our security properties into their website, as it is the de facto source of truth for informing users about the risks of rollups. Additionally, we have collaborated closely with the Matter Labs team to help them design secure, verified mechanisms for contracts operating the zkSync Era rollup.
How we can use Alloy models. Alloy models can provide high-level specification and guidance for the low-level implementation of the described protocols. Another promising direction is generating tests from Alloy models to identify corner cases [24] and validate Solidity contracts against these scenarios. It is worth mentioning that related tools like TLA+ [25] and Quint [26] offer closer connections to code and facilitate code extraction.
For more detailed examples and potential extensions of our model, refer to Appendix B.
@@ 6 Discussion
Simple 295 1,203,661 9.099
Forced Queue 529 1,996,273 136.104
Blacklist 721 2,200,981 399.724
Upgradeability 970 2,468,911 388.620
@@
Figure 4: Various statistics for each rollup mechanism. The last column indicating the solving time uses a scope of 5.
Figure 5: Log-scale graph of analysis time for increasing scopes for our most complex model, i.e. the model for upgradeability.
7 Measurements
@@
In this paper, we use Alloy Analyze  6 to verify the invariants and assertions of various L2 mechanisms. This involved adapting our model for each ZK Rollup mechanism and then systematically testing them within the Alloy environment. The scope for our tests was set up to allow us to specify more fine-grained bounds on each type but does not imply any limitation on the overall number of system states that can be explored. All experiments were performed on an Intel Core i7-9750H CPU with 16 GB of RAM.
@@
For each case, we quantified the number of new lines of Alloy code written, the total number of clauses generated by the analyzer, and the time to solve these clauses using an SAT solver. Figure 4 shows that the complexity and the time required to solve the clauses increase with more complex rollup mechanisms. Figure 5 depicts the running time for different scope values for our most complex model, demonstrating a significant increase in the SAT solver time as the scope increases. The observed exponential growth is typical of SAT-based analyses.
8 Related Work
Formal Methods for Web Security. Our work is inspired by the seminal work of Akhawe et al. [17] on applying formal modeling to analyze the security of web mechanisms and applications. Specifically, this work provides a formal model of the web platform to dissect the security of various web mechanism designs using Alloy. Their analysis included the specification of threat models tailored for evaluating the security of web platforms. Analogously, we crafted our analysis around the strongest conceivable adversary (i.e., a malicious L2), examining rollup designs under this threat model. Moreover, Akhawe et al. [17] established a set of broadly applicable security goals and analyzed various security properties for web applications. Similarly, we defined fundamental security goals tailored for rollup users, defined security properties, and systematically evaluated different scenarios using Alloy. For those interested in formal methods for security, the recent survey by Kulik et al. [27] offers a detailed examination of formal methods accessible to designers of security-critical systems.
Formal methods in blockchain and ZKPs. There is a growing body of work on verifying different aspects of blockchain, from smart contacts to consensus algorithms and ZK proof generation. We mention only some relevant work here. For smart contract verification, existing research endeavors primarily aim to either demonstrate the absence of common vulnerabilities or rely on user-defined specifications to verify correctness. Notable surveys by Murray and Anisi [28] and Garfatta et al. [29] offer comprehensive overviews of formal verification methods tailored for Solidity smart contracts. However, it is important to note that our work diverges from this trajectory, focusing instead on providing assurances regarding the design of mechanisms implemented in smart contracts that operate rollups.
On the front of formal verification for Byzantine fault tolerance (BFT) in blockchain systems, seminal works by Tholoniat and Gramoli [30] and Yoo et al. [31] have contributed significantly. Regarding ZKPs, recent efforts have explored the verification of verifiers for ZK-rollups [20]. Additionally, several initiatives have focused on formally proving the absence of common bugs in ZK circuits [18]. In contrast to those efforts, we focus on the design of mechanisms that are complementary to ZKPs.
Rollup Security. Prior works have focused on qualitatively analyzing optimistic and ZK-Rollups. Gorzny et al. [32] established a wishlist of properties that an escape hatch mechanism (also known as a forced queue) should possess to be considered trustworthy. Their paper focuses on qualitative characteristics that escape hatches should exhibit, such as executing arbitrary transactions, being supported by default in a rollup, and efficiency. In contrast, our paper formally defines the security properties of escape hatches and analyzes various scenarios using Alloy.
In their recent work, Koegl et al. [33] presented nine main attack threats for rollups, briefly describing potential attacks in both optimistic and ZK-Rollups, such as censorship attacks, DoS attacks, and client vulnerability risks. Our work extends some of these potential attacks by formally modeling ZK-Rollups, presenting a clear and strong adversary, and outlining the security properties necessary to ensure a secure ZK Rollup design.
Finally, complementing our work, Gorzny and Derka [34] presented a framework to qualitatively evaluate existing rollups based on different dimensions: familiarity, finality time, modularity, and maturity. In comparison to the aforementioned works, we are the first to formally specify the main properties of ZK-Rollups and analyze various designs using Alloy to provide formal guarantees of the security of specific ZK rollup designs.
9 Conclusion
Rollups are heralded as the foremost solution to blockchain scalability, collectively securing over \$40 billion in value. However, recent incidents have led to censorship and halting block production on rollups. Despite these astronomical monetary amounts, the security properties of rollups remain inadequately understood.
Our work introduces an extensible formal model for ZK-Rollups that facilitates not only the design and evaluation of rollups’ critical mechanisms but also enhances the practical understanding of their security properties. Using Alloy, a language that has been used for modeling complex systems for more than a decade, we create models that are accessible to developers, security auditors, and researchers alike, enabling automatic reasoning with rapid results. By making our models public, we encourage their use for evaluating concrete implementations. Finally, the comprehensive security properties described in our work could be showcased on platforms dedicated to rollup security, like L2Beat. This would give developers and users a clear and tangible understanding of the security assurances each rollup provides. Our primary objective is to lay the foundation for the creation and advancement of secure L2 solutions that are not only informally recognized but also verifiably proven.
Acknowledgements
We thank Franck Cassez, Bartek Kiepuszewski, Piotr Szlachciak, Anton Astafiev, Marco Stronati, Antonio Locascio, Kalman Lajko, and Matej Pavlovic for their insightful feedback.
Appendix A Alloy Metamodel for the Forced Queue and Upgradeability Model
Figure 6 illustrates the metamodel used for formalizing the forced queue and upgradeability model in Alloy. The metamodel depicts various components and their connections within the system.
Figure 6: The metamodel of our formalization of the forced queue and upgradeability in Alloy.
@@ Appendix B Model Extensions
Analyzing Withdrawal Windows in Optimistic Rollups. Optimistic rollups that have been deployed generally employ a seven-day withdrawal period. Although numerous reasons exist for the necessity of this duration [35], newer methods propose shortening these periods. Nevertheless, the potential security compromises of these more assertive suggestions [36] still require thorough examination.

### v3

@@ 5.1. Design-time Testing
Two issues could arise from merely enforcing the upgrade after a set period. For the first example, consider the case where we use a forced queue and upgradeability described in Section 4.5. However, we deploy the upgrade based solely on the timeout (i.e., not taking care of locking and emptying the forced queue before deploying the upgrade). Figure 3 demonstrates the issue that might arise in that case produced by Alloy. Specifically, what happens is that first, we start with an empty forced queue, and then a user adds an input to the queue (Input_1). Next, an announcement occurs, and then the timeout of that announcement is reached. This enforces the upgrade, updating the blacklist to include Input_1. The result is that L2 becomes frozen because the head (Input_1) of the queue is blacklisted. Another problem could occur if the upgrade involves changing the functionality of the rollup and a user might submit a transaction to be forced before the upgrade because they disagree with the change and want to exit before the upgrade. However, because other transactions are in front of it (whether benignly or maliciously added), their transaction will not be processed. Additionally, the L2 might censor the user’s transactions, preventing the user from exiting before the upgrade.
These examples demonstrate severe issues with existing designs. Notably, such an issue existed in the zkSync Era’s design (not in production, as this mechanism has not been deployed yet), and we have worked closely with them to adapt their design to fix the issue. Our model presented in Section 4 is immune to the described attacks. Finally, in Appendix A, we provide detailed statistics on our alloy model and a detailed analysis of the performance and scalability of our model.
rule guaranteedProcessing(env e) {
require forcedQueue.length > 0;
uint256 tx == forcedQueue[0];
finalizeState(e);
assert !forcedQueue.contains(tx);
(a) Certora rule encoding FQP1 for formal verification.
function testGuaranteedProcessing() public {
vm.assume(forcedQueue.length > 0);
uint256 tx == forcedQueue[0];
finalizeState();
assert(!forcedQueue.contains(tx));
(b) Foundry (Forge) test encoding FQP1 as a property-based invariant.
Figure 4. Example of using Certora (a) and Foundry (b) to express the Guaranteed Processing property (FQP1).
5.2. Code-level Verification/Testing
In the previous section, we demonstrated how our formal model aids in the high-level design verification of ZK-Rollup mechanisms. Here, we propose an extension of this approach to the implementation level (smart contracts), illustrating how the properties defined in Section 4, i.e., Forced Queue Properties (FQPs), Blacklisting Properties (BPs), and Upgradeability Properties (UPs), can be employed to test actual smart contract implementations of these mechanisms.
Formal verification and property-based testing of smart contracts. Formal verification tools like Certora (Nandi et al., 2025) and property-based testing frameworks such as Foundry (Foundry contributors, 2021) and Echidna (Grieco et al., 2020) are widely adopted in the DeFi ecosystem, with protocols like Lido and AAVE utilizing them to ensure contract correctness (Chaliasos et al., 2024a). Certora allows developers to write specifications in the Certora Verification Language (CVL), enabling the formal verification of smart contracts against defined properties. Foundry and Echidna facilitate property-based testing by generating a wide range of inputs to test contract behaviors against specified invariants.
End-to-end analysis of ZK-Rollups. Our methodology involves adapting the formally verified properties from our Alloy model (FQPs, BPs, and UPs) to the smart contracts of a selected rollup. These adapted properties are then verified/tested using frameworks like Certora or Foundry. Given that Alloy has formally verified these properties, any counterexamples found by the employed tool indicate implementation bugs. This approach ensures the implementation aligns with the formally verified design, bridging the gap between high-level specifications and practical deployment.
To illustrate, consider FQP1 (Section 4.3). This property ensures that if the forced queue is non-empty and a new state is finalized, then the head of the forced queue must be processed and removed. For simplicity, consider a smart contract of a rollup with a queue forcedQueue containing transaction IDs. Figure 4(a) depicts the rule for Certora, while Figure 4(b) shows the invariant for Foundry. This example demonstrates how Alloy properties can be translated into formal verification rules or invariants, ensuring that implementations adhere to their specifications. In Appendix B, we explore more potential future work of automating the whole process.
Analyze Scroll’s forced queue. Scroll implements only the forced queue mechanism from the mechanisms presented in this paper. Unlike our general model, Scroll’s implementation enforces the forced queue only when a timeout occurs, i.e., when the head of the queue has been delayed beyond a predefined threshold. At that point, the rollup cannot progress until it processes the timed-out forced queue messages. To accurately capture this behavior, we extended our Alloy model by refining the forced queue properties (FQP) to encode timeouts and introducing four Scroll-specific properties: (i) Rolling hash integrity to ensure correct hashing of queued messages, (ii) Enforced mode activation to verify that enforced mode is triggered after a timeout, (iii) Mode consistency to guarantee mutual exclusivity between enforced and normal modes, and (iv) Fee payment to confirm that forced queue messages cover their required fees.
We translated this adapted Alloy model into a property-based testing suite using Foundry, implementing all specified properties in approximately 1,000 lines of Solidity code. Each property was encoded as an invariant checked after random operations, with a total of 10,000 runs. No violations were detected in the forced queue implementation of Scroll. The full adaptation and testing process took three days, aided significantly by Scroll’s existing support for Foundry-based unit tests and mock operations such as batch proving and verification. The existing infrastructure was a key reason for selecting Foundry over Certora.
6. Discussion
Alloy’s Bounded Scope and Security Guarantees. Alloy exhaustively explores all behaviours within a bound on object count and trace length; success inside that window does not prove universal correctness (Brunel et al., 2021). It is therefore a powerful bug-finder, but not a complete verifier. We ran the analyser with at most 10 objects (scope) and 10-step traces. This setting follows the small-scope hypothesis, which holds that most design errors admit small counter-examples (Jackson and Damon, 1996; Jackson, 2019). The flaw in Fig. 3 arises with a single forced transaction and one upgrade event; adding a second queued transaction reveals the subtler ‘blocked-exit’ scenario. To gauge this approach, we injected five synthetic faults: four surfaced at scope $\leq 5$, the fifth at scope 8. This suggests a scope of 8 would suffice, yet we conservatively used up to 10. Appendix A shows how SAT-solver time grows exponentially with scope, reinforcing the need to cap it. Importantly, our model may miss issues that necessitate a larger scope to be manifested.
Analysing Further L2 Mechanisms. Our model focuses on mechanisms enforced by L1 contracts: forced queues, blacklisting, and upgradeability, because these must remain correct regardless of how the L2 is engineered; they are essential for a rollup to inherit the security properties of its base chain. Other mechanisms that touch multiple layers have not been analyzed. Withdrawals, for instance, interact with both layers: an exit begins on L2 but is finalised on L1 only after the corresponding L2 transaction is settled, so modelling them faithfully would require incorporating the bridge contract and relevant L2 logic, an extension we leave to future work that can build on prior bridge analyses (Lee et al., 2023). Further, data-availability guarantees, which depend on blob-space economics or external DA layers, introduce complexities beyond our present scope. In short, we prioritise L1-relevant mechanisms that must remain sound even under a hostile L2; extending the model to cross-layer withdrawals is a natural next step, while DA and sequencer-level protocols will likely require different analyses.
Deployment Status and Practical Relevance of ZK-Rollups. ZK-Rollups are no longer research prototypes: according to L2Beat (July 2025), six of the ten largest rollups by TVL are ZK-Rollups, together securing well over \$3 billion in user assets [^15] https://l2beat.com/scaling/summary. The remaining four are optimistic rollups, yet even optimistic rollups could potentially move towards ZK proofs. For example, OP-Succinct (Succinct Labs) aims to let existing optimistic stacks switch to validity proofs with minimal code changes.  [^16] https://github.com/succinctlabs/op-succinct Of the three mechanisms we formalize, forced exits and timelocked upgrades are already on every major rollup’s public roadmap; nevertheless, only Scroll has a deployed forced-queue today (c.f., Section 5.2). The safe-blacklisting mechanism is less ubiquitous but increasingly demanded by institutionally focused chains that require compliance; our analysis shows that simple ‘on-the-spot’ blacklists could break liveness (Section 5.1), while our model preserves user guarantees. In conclusion, ZK-Rollups are widely deployed, the proposed mechanisms are on the critical path for all of them, and the presented formal model has already proven its practical value by uncovering potential issues in the interactions between mechanisms and helping test critical properties of a live system.
7. Related Work
Formal Methods for Web Security. Our work is inspired by the seminal work of Akhawe et al. (Akhawe et al., 2010) on applying formal modeling to analyze the security of web mechanisms and applications. Specifically, this work provides a formal model of the web platform to dissect the security of various web mechanism designs using Alloy. Their analysis included the specification of threat models tailored for evaluating the security of web platforms. Analogously, we crafted our analysis around the strongest conceivable adversary (i.e., a malicious L2), examining rollup designs under this threat model. Moreover, Akhawe et al. (Akhawe et al., 2010) established a set of broadly applicable security goals and analyzed various security properties for web applications. Similarly, we defined fundamental security goals tailored for rollup users, defined security properties, and systematically evaluated different scenarios using Alloy. For those interested in formal methods for security, the recent survey by Kulik et al. (Kulik et al., 2022) offers a detailed examination of formal methods accessible to designers of security-critical systems.
Formal methods in blockchain and ZKPs. There is a growing body of work verifying different aspects of blockchain, from smart contracts to consensus algorithms and ZK proof generation. We mention only some relevant work here. For smart contract verification, existing research endeavors primarily aim to demonstrate the absence of common vulnerabilities or rely on user-defined specifications to verify correctness. Notable surveys by Murray and Anisi (Murray and Anisi, 2019) and Garfatta et al. (Garfatta et al., 2021) offer comprehensive overviews of formal verification methods tailored for smart contracts. However, it is important to note that our work diverges from this trajectory, focusing instead on providing assurances regarding the design of mechanisms implemented in smart contracts that operate rollups.
On the front of formal verification for Byzantine fault tolerance (BFT) in blockchain systems, seminal works by Tholoniat and Gramoli (Tholoniat and Gramoli, 2022) and Yoo et al. (Yoo et al., 2019) have contributed significantly. Regarding ZKPs, recent efforts have explored the verification of verifiers for ZK-rollups (Firsov and Livshits, 2024). Additionally, several initiatives have focused on formally proving the absence of common bugs in ZK circuits (Chaliasos et al., 2024b). In contrast to those efforts, we focus on the design of mechanisms that are complementary to ZKPs.
Rollup Security. Prior works have focused on qualitatively analyzing optimistic and ZK-Rollups. Gorzny et al. (Gorzny et al., 2022) established a wishlist of properties that an escape hatch mechanism (also known as a forced queue) should possess to be considered trustworthy. Their paper focuses on qualitative characteristics that escape hatches should exhibit, such as executing arbitrary transactions, being supported by default in a rollup, and efficiency. In contrast, our paper formally defines the security properties of escape hatches and analyzes various scenarios using Alloy, enabling the identification of vulnerabilities in flawed designs.
In their recent work, Koegl et al. (Koegl et al., 2023) presented nine main attack threats for rollups, briefly describing potential attacks in both optimistic and ZK-Rollups, such as censorship attacks, DoS attacks, and client vulnerability risks. Our work extends some of these potential attacks by formally modeling ZK-Rollups, presenting a clear and strong adversary, and outlining the security properties necessary to ensure a secure ZK Rollup design. Finally, complementing our work, Gorzny and Derka (Gorzny and Derka, 2024) presented a framework to qualitatively evaluate existing rollups based on different dimensions: familiarity, finality time, modularity, and maturity.
In comparison to the aforementioned works, we are the first to formally specify the main properties of ZK-Rollups and analyze various designs using Alloy, providing formal guarantees of the security of specific ZK rollup designs. In essence, while prior works provide generic qualitative directions, we have developed the first tooling based on formal methods to help practitioners identify potential issues in the security-critical mechanisms of rollups.
8. Conclusion
Rollups are heralded as the foremost solution to blockchain scalability, collectively securing over \$30 billion in value. However, recent incidents have led to censorship and halting block production on rollups. Despite these astronomical monetary amounts, the security properties of rollups remain inadequately understood.
Our work introduces an extensible formal model for ZK-Rollups that facilitates the design and evaluation of rollups’ critical mechanisms and enhances the understanding of their security properties. Using Alloy, a language used for modeling complex systems for over a decade, we create models accessible to developers, security auditors, and researchers, enabling automatic reasoning with rapid results. We further show how the security properties defined in Alloy can be used to verify or test the smart contracts implementing rollup logic on-chain, using formal verification or property-based testing. Our primary objective is to lay the foundation for creating and advancing secure L2 solutions that are verifiably proven.
Acknowledgements.
We would like to thank the anonymous CCS reviewers for their insightful feedback on previous versions of the paper. D. Firsov was supported by the Estonian Research Council grant no. PSG749.
Appendix A Alloy Measurements
@@
Simple 295 1,814,483 63.603
Forced 529 3,633,032 89.623
Blacklist 721 5,461,695 152.688
Upgrade 970 6,149,438 125.703
Upgrade+Blacklist - 7,978,101 188.768
@@
Figure 5. Various cumulative statistics for each rollup mechanism. The No. of Clauses and Solve time are for scope 5 and 10 steps.
Figure 6. Log-scale graph of analysis time for 5 and 10 scopes. The colored part depicts the variance across properties.
@@
In this paper, we use Alloy Analyzer 6 to verify the invariants and assertions of various L2 mechanisms. This involved adapting our model for each ZK-Rollup mechanism and then systematically testing them within the Alloy environment. The scope for our tests was set up to allow us to specify more fine-grained bounds on each type, but it does not imply any limitation on the overall number of system states that can be explored. All experiments were performed on a MacBook Pro Apple M1 Max CPU with  32 GB of RAM.
@@
For each case, we quantified the number of new lines of Alloy code written, the total number of clauses generated by the analyzer, and the time to solve these clauses using an SAT solver. Figure 5 shows the lines of code required for each mechanism, the cumulative number of clauses of the properties of each mechanism, and the cumulative solve time in seconds for scope 5 and steps 1–10. The complexity and the time required to solve the clauses increase with more complex rollup mechanisms. Figure 6 depicts the running time for different step values for scopes 5 and 10, demonstrating a significant increase in the SAT solver time as the scope increases, as well as the steps increase. The observed exponential growth is typical of SAT-based analyses. Finally, the colored area demonstrates the variance between checking different properties.
@@ Appendix B Model Extensions
Analyzing Withdrawal Windows in Optimistic Rollups. Optimistic rollups that have been deployed generally employ a seven-day withdrawal period. Although numerous reasons exist for the necessity of this duration (Fichter, [n. d.]), newer methods propose shortening these periods. Nevertheless, the potential security compromises of these more assertive suggestions (Zeeve, [n. d.]) still require thorough examination.